Md:

Juanita and Brett had $280 altogether. After Juanita spent 1/2 of her money and Brett spent 1/4 of his money, they each had the same amount of money left. How much money did Juanita have at first?

Hero:

"Juanita and Brett had $280 altogether" translates to: j + b = 280 "After Juanita spent 1/2 of her money and Brett spent 1/4 of his money, they each had the same amount of money left" translates to j - (1/2)j = b - (1/4)b

Hero:

So what we have is a system where j + b = 280 j - j/2 = b - b/4 and j represents the amount of money juanita had originally and b represents the amount of money brett had originally and we are asked to solve for j.

Hero:

Are you following @Md ?

Ultrilliam:

and hes gone...

Ultrilliam:

And hes back

Ultrilliam:

@Md